# Kernel PCA for Out-of-Distribution Detection

### v1

@@ 4.1 Comparisons with Nearest Neighbor Searching
The comparisons with KNN [7] cover both the benchmarks. Following the setups in KNN, for fair comparisons, we evaluate models trained via the standard cross-entropy loss and models trained via the supervised contrastive learning [40], and adopt the same checkpoints released by KNN: ResNet18 [41] on CIFAR10 and ResNet50 on ImageNet-1K. Here, the scoring function of CoP and CoRP is $S(\boldsymbol{\hat{x}})=-e^{\Phi}(\boldsymbol{\hat{x}})$.
Tab.1 presents empirical results on the ImageNet-1K benchmark. In standard training, our CoRP shows superior detection performance over KNN with lower FPR and higher AUROC values averaged over OoD data sets. In supervised contrastive learning, both CoP and CoRP outperform other baseline results on each OoD data set. These results show that the proposed KPCA exploring non-linear patterns is more advantageous than the nearest neighbor searching and all compared methods. Besides, the further improvements of CoRP over CoP also verify the effectiveness of the distance-preserving property of Gaussian kernel $k_{\rm gau}$ on top of cosine kernel $k_{\rm cos}$ for OoD detection.
In addition, KPCA outperforms KNN on the CIFAR10 benchmark with improved OoD detection performances, and also shows a much cheaper time expense in inference, which we leave to the Appendix for more details.
Table 2: Comparisons on the detection performance between the regularized reconstruction error [8] and the proposed CoP and CoRP fused with other OoD scores (MSP, Energy, ReAct and BATS) on each OoD data set (ResNet50 pre-trained on ImageNet-1K).
method OoD data sets
iNaturalist SUN Places Textures AVERAGE
@@ 4.2 Comparisons with Regularized Reconstruction Errors
In [8], the authors proposed a regularized reconstruction error and a fusion strategy to combine with existing OoD scores. The regularized reconstruction error $e_{\rm reg}(\boldsymbol{\hat{z}})$ is calculated in the original $\boldsymbol{z}$-space with the PCA reconstruction error $e(\boldsymbol{\hat{z}})$ Eq.(4) and the normalization scale $\left\|\boldsymbol{\hat{z}}\right\|_{2}$ by : $e_{\rm reg}(\boldsymbol{\hat{z}})=\frac{\|\boldsymbol{U}_{q}\boldsymbol{U}_{q}^{\top}(\boldsymbol{\hat{z}}-\boldsymbol{\mu})-(\boldsymbol{\hat{z}}-\boldsymbol{\mu})\|_{2}}{\left\|\boldsymbol{\hat{z}}\right\|_{2}}.$ The authors claim a fusion strategy for $e_{\rm reg}$ and other existing OoD scores, e.g., the Energy [5], as $(1-e_{\rm reg}(\boldsymbol{\hat{z}}))\cdot S_{\rm Energy}(\boldsymbol{\hat{z}})$ with a strong OoD detector. Following the settings in [8], for a fair comparison, the proposed KPCA-based reconstruction error $e^{\Phi}$ is also evaluated under the same framework. The fused OoD scores include MSP [11], Energy [5], ReAct [14] and BATS [15]. The detection experiments are executed on the ImageNet-1K benchmark with pre-trained ResNet50 and MobileNet [42] checkpoints from PyTorch [43].
@@
Tab.2 presents the comparisons between [8] and ours on ResNet50. In almost all the cases, both the KPCA-based CoP and CoRP outperform the regularized reconstruction error [8] on all the OoD data sets with substantially improved FPR and AUROC values. Specifically, when fused with the ReAct method [14], the CoRP achieves new SOTA OoD detection performance among various detectors. All these experiment results indicate that an appropriately mapped $\Phi(\boldsymbol{z})$-space benefits the OoD detection than the original $\boldsymbol{z}$-space as the non-linearity in $\boldsymbol{z}$-space gets alleviated by the feature mapping $\Phi$. Our work provides 2 selections for $\Phi$ and empirically verifies their superior effectiveness. See comparisons on MobileNet in the Appendix.
Figure 2: A sensitivity analysis on the explained variance ratio in CoP. The average detection FPR and AUROC values of OoD data sets in CIFAR10 and ImageNet-1K benchmarks are reported.
4.3 Sensitivity Analysis
This section provides sensitivity analyses to show the effects of hyper-parameters in CoP and CoRP. The number of columns (components) of the dimensionality-reduction matrix $\boldsymbol{U}_{q}^{\Phi}$, i.e., $q$, is significant for the detection performance, as it determines how much information captured by the subspace, where the InD and OoD data is projected onto. The selection of $q$ is determined by the explained variance ratio, reflecting the informativeness of the kept principal components. Fig.2 illustrates the detection performance of CoP under varied explained variance ratios. On CIFAR10 and ImageNet-1K benchmarks, a mild value of the explained variance ratio is suggested with around 90% for keeping the components. More detailed sensitivity analyses can refer to the Appendix on CoRP with the bandwidth $\gamma$ in the Gaussian kernel and the dimension $M$ of RFFs.
Figure 3: Ablation studies on the cosine feature mapping $\phi_{\rm cos}$ in CoP and CoRP. The detection FPR values of each OoD data set in CIFAR10 and ImageNet-1K benchmarks are reported.
4.4 Ablation Studies
Ablation studies on the cosine feature mapping $\phi_{\rm cos}$ are conducted to illustrate the significance in benefiting the proposed KPCA OoD detector. Specifically, CoP without $\phi_{\rm cos}$ reduces to PCA on the $\boldsymbol{z}$-space, and CoRP without $\phi_{\rm cos}$ reduces to PCA on the mapped $\phi_{\rm RFF}(\boldsymbol{z})$-space w.r.t. a Gaussian kernel. Fig.3 shows the corresponding detection FPR values on each OoD data set in CIFAR10 and ImageNet-1K benchmarks.
As illustrated in Fig.3, the cosine feature mapping $\phi_{\rm cos}$ serves as an essential basis for the effectiveness of PCA in the mapped feature space for OoD detection, shown by the substantial increased FPR values of CoP and CoRP without $\phi_{\rm cos}$. Such ablation studies are also in line with the findings in KNN [7], where the $\ell_{2}$-normalization on penultimate features $\boldsymbol{z}$ is critical for the nearest neighbor searching.
5 Analytical Discussions
@@
In CoP and CoRP, KPCA is executed with the covariance matrix of mapped features $\Phi(\boldsymbol{z})$. In contrast, in the classic KPCA [10], such feature mappings $\Phi$ are not explicitly given, and it rather works with a kernel function applied to $\boldsymbol{z}$. In this section, we supplement our covariance-based KPCA with its kernel function implementation, including theoretical discussions and comparison experiments on OoD detection. Our CoP and CoRP are shown to be more effective and efficient than their counterparts that employ kernel functions.
@@
In the classic KPCA, the kernel trick enables projections to the principal subspace via kernel functions without calculating $\Phi$. However, how to map the projections in the principal subspace back to the original $\boldsymbol{z}$-space remains a non-trivial issue, known as the pre-image problem [44], which makes it problematic to calculate reconstructed features via kernel functions. To address this issue, Proposition 1 shows a flexible way to directly calculate reconstruction errors without building reconstructed features, so as to apply the kernel trick, shown in Proposition 2.
@@
Proposition 1 implies that the reconstruction error equals to the norm of projections in the residual $\boldsymbol{U}^{\Phi}_{p}$-subspace, i.e., the subspace consisting of those principal components that are not kept, see proofs in the Appendix. Accordingly, as typically done in the classic KPCA, we can introduce a kernel function to perform dimension reduction, but to the residual subspace, and then calculate the norms of the reduced features as reconstruction errors, illustrated by Proposition 2.
Given a kernel function $k(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, we have a kernel matrix $\boldsymbol{K}\in\mathbb{R}^{N_{\rm tr}\times N_{\rm tr}}$ on training data with $\boldsymbol{K}_{i,j}=k(\boldsymbol{z}_{i},\boldsymbol{z}_{j})$, and a vector $\boldsymbol{k}_{\boldsymbol{\hat{z}}}\in\mathbb{R}^{N_{\rm tr}}$ with the $i$-th element $k(\boldsymbol{z}_{i},\boldsymbol{\hat{z}})$ for a new sample $\boldsymbol{\hat{z}}$. We have Proposition 2.
@@
The KPCA reconstruction error $e^{k}(\boldsymbol{\hat{z}})$ w.r.t. a kernel function $k$ can be calculated as follows:
@@
where $\boldsymbol{A}\in\mathbb{R}^{N_{\rm tr}\times l}$ includes $l$ eigenvectors of the kernel matrix $\boldsymbol{K}$ w.r.t. the top-$l$ smallest eigenvalues.
According to Proposition 2, now CoP and CoRP can be implemented via kernel functions. For CoP, we just directly apply the cosine kernel $k_{\rm cos}$ in Eq.(9). For CoRP, we should adopt the Gaussian kernel $k_{\rm gau}$ on the $\ell_{2}$-normalized inputs $\frac{\boldsymbol{z}}{\left\|\boldsymbol{z}\right\|_{2}}$. Fig.4 shows comparisons on the OoD detection performance based on reconstruction errors calculated by CoP/CoRP and their kernel function implementations.
In Fig.4, the detection performance of KPCA with kernel functions is evaluated by varying the explained variance ratio of the kernel matrix $\boldsymbol{K}$. The larger the explained variance ratio, the smaller the dimension $l$ of $\boldsymbol{A}$. The best detection results achieved by CoP/CoRP are illustrated. Clearly, regrading the OoD detection performance, reconstruction errors $e^{k}$ calculated by kernel functions are less effective than those calculated explicitly in the mapped $\Phi(\boldsymbol{z})$-space.
@@
This section supplements our covariance-based KPCA with its counterpart that adopts kernel functions. A flexible way to calculate reconstruction errors via kernel functions for OoD detection is presented. Even though, the KPCA with kernel functions is less efficient than CoP/CoRP in two aspects: (i) The time expense of eigendecomposition on the kernel matrix by the former is more expensive than that on the covariance matrix by the latter, since the dimension of the kernel matrix is much larger than that of the covariance matrix. (ii) The former requires an $\mathcal{O}(N_{\rm tr})$ complexity of calculating $\boldsymbol{k}_{\boldsymbol{\hat{z}}}$ in inference, much higher than the $\mathcal{O}(1)$ complexity of the latter.
@@ 6 Conclusion
As PCA reconstruction errors fail to distinguish OoD data from InD data on the penultimate features $\boldsymbol{z}$ of DNNs, kernel PCA is introduced for its non-linearity in the manner of employing explicit feature mappings. To find an appropriate feature mapping that can characterize the non-linear patterns in InD and OoD features, we take a kernel perspective on an existing KNN detector [7], and thus propose two explicit feature mappings w.r.t. a cosine kernel and a cosine-Gaussian kernel. The mapped $\Phi(\boldsymbol{z})$-space enables PCA to extract principal components that well separate InD and OoD data, leading to distinguishable reconstruction errors. Extensive empirical results have verified the improved effectiveness and efficiency of the proposed KPCA with new SOTA OoD detection performance. Finally, theoretical discussions and associated experiments are provided to bridge the relationships between our covariance-based KPCA and its kernel function implementation.
@@
One limitation of the KPCA detector is that the two specific kernels are still manually selected with carefully-tuned parameters. It remains a valuable topic in the OoD detection task whether the parameters of kernels could be learned from data according to some optimization objective, i.e., kernel learning. We hope that the proposed two effective kernels verified empirically in our work could benefit the research community as a solid example for future studies.
Supplementary Material for
Kernel PCA for Out-of-Distribution Detection
I Details of OoD Detectors Involved in Comparison Experiments
@@
In this section, we elaborate the scoring function $S(\cdot)$ of the OoD detectors included in the comparison experiments of Sec.4.1 and Sec.4.2. Given a well-trained DNN $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$ with inputs $\boldsymbol{x}\in\mathbb{R}^{d}$, the outputs are $c$-dimension logits $f(\boldsymbol{x})\in\mathbb{R}^{c}$ w.r.t. $c$ classes. The DNN $f$ learns features $\boldsymbol{z}\in\mathbb{R}^{m}$ of $\boldsymbol{x}$ before the last linear layer, i.e., the penultimate features $\boldsymbol{z}$.
@@
MSP [11] employs the softmax function on the output logits and takes the maximum probability as the scoring function. Given a test sample $\boldsymbol{\hat{x}}\in\mathbb{R}^{d}$, its MSP score is
$S_{\rm MSP}(\boldsymbol{\hat{x}})=\max\left({\rm softmax}(f(\boldsymbol{\hat{x}}))\right).$ (I.1)
@@
$S_{\rm ODIN}(\boldsymbol{\hat{x}})=\max\left({\rm softmax}(\frac{f(\boldsymbol{\hat{x}}_{\rm a})}{T})\right),$ (I.2)
@@
Mahalanobis [16] employs the Mahalanobis score to perform OoD detection. The DNN outputs at different layers are modeled as a mixture of multivariate Gaussian distributions and the Mahalanobis distance is calculated. Then, a linear regressor is trained to achieve a weighted Mahalanobis distance at different layers as the final detection score. To train the linear regressor, the training data and the corresponding adversarial examples are adopted as positive and negative samples, respectively.
$S^{l}_{\rm mahal}(\boldsymbol{\hat{x}})$ $=\max_{i}\ -(f^{l}(\boldsymbol{\hat{x}})-\boldsymbol{\mu}^{l}_{i})^{T}\boldsymbol{\Sigma}_{l}(f^{l}(\boldsymbol{\hat{x}})-\boldsymbol{\mu}_{i}^{l}),$ (I.3)
@@
Energy [5] adopts an energy function on logits since energy is better aligned with the input probability density:
$S_{\rm energy}(\boldsymbol{\hat{x}})=\log\sum_{i=1}^{c}\exp(f_{i}(\boldsymbol{\hat{x}})),$ (I.4)
where $f_{i}(\boldsymbol{\hat{x}})$ denotes the $i$-th element in the $c$-dimension output logits.
@@
$\boldsymbol{\bar{z}}=\min\left\{\boldsymbol{z},\alpha\right\},$ (I.5)
where $\alpha$ is a pre-defined constant. The clipped features $\boldsymbol{\bar{z}}$ then pass through the last linear layer and yield modified logits. Other logits-based OoD methods such as Energy could be applied on the modified logits to produce detection scores.
KNN [7] is a simple but time-consuming and memory-inefficient detector since it performs nearest neighbor search on the $\ell_{2}$-normalized penultimate features between the test sample and all the training samples. The negative of the ($k$-th) shortest $\ell_{2}$ distance is set as the score for a new sample $\boldsymbol{\hat{x}}$:
$S_{\rm knn}(\boldsymbol{\hat{x}})=-\min_{i:1,\cdots,N_{\rm tr}}\left\|\frac{\boldsymbol{\hat{z}}}{\|\boldsymbol{\hat{z}}\|_{2}}-\frac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|_{2}}\right\|_{2},$ (I.6)
where $\boldsymbol{z}_{i}$ denotes the penultimate features of the $i$-th training sample in the training set of a size $N_{\rm tr}$. The key of this detector is the $\ell_{2}$-normalization on the features.
@@
$f_{\rm DICE}(\boldsymbol{\hat{x}})\triangleq\left(\boldsymbol{M}\odot\boldsymbol{W}\right)^{\top}\boldsymbol{\hat{z}}+\boldsymbol{b}.$ (I.7)
$\odot$ is the element-wise multiplication, and $\boldsymbol{M}\in\mathbb{R}^{m\times c}$ is a masking matrix whose elements are determined by the element-wise multiplication between the $i$-th column $\boldsymbol{w}_{i}\in\mathbb{R}^{m}$ in $\boldsymbol{W}$ and the penultimate features $\boldsymbol{\hat{z}}$: $\boldsymbol{w}_{i}\odot\boldsymbol{\hat{z}}$. Then, similar as ReAct, logits-based detectors could be executed on the modified logits $f_{\rm DICE}(\boldsymbol{\hat{x}})$ to produce detection scores.

### v3

@@ 5.1 Comparisons with nearest neighbor searching
The comparisons with KNN [7] cover both the benchmarks. Following the setups in KNN, for fair comparisons, we evaluate models trained via the standard cross-entropy loss and models trained via the supervised contrastive learning [49], and adopt the same checkpoints released by KNN: ResNet18 [50] on CIFAR10 and ResNet50 on ImageNet-1K. Here, the scoring function of CoP and CoRP is $S(\boldsymbol{\hat{x}})=-e^{\Phi}(\boldsymbol{\hat{x}})$ in Algorithm 1.
Table 1 presents empirical results of ResNet50 on the ImageNet-1K benchmark. In the standard training, our CoRP shows superior detection performance over KNN with lower FPR and higher AUROC values averaged over multiple OoD data sets. In supervised contrastive learning, both CoP and CoRP outperform other baseline results on each OoD data set. These results show that the proposed KPCA exploring non-linear patterns is more advantageous than the nearest neighbor searching and all compared methods. Besides, the further improvements of CoRP over CoP also verify the effectiveness of the distance-preserving property of the Gaussian kernel $k_{\rm gau}$ on top of the cosine kernel $k_{\rm cos}$ for OoD detection.
On the other hand, regarding the computational complexity in inference, Table 2 empirically shows the superior $\mathcal{O}(1)$/${\cal O}(M)$ time and memory complexity of CoP/CoRP over the $\mathcal{O}(N_{\rm tr})$ of KNN, including: (i) the inference time of the nearest neighbor search in KNN and the reconstruction error calculation in CoP/CoRP; (ii) the storage of the InD training features in KNN and the $\boldsymbol{U}^{\Phi}_{q}$ and $\boldsymbol{\mu}^{\Phi}$ in CoP/CoRP. To be specific, for KNN, storing and iterating all the $N_{\rm tr}=1,281,167$ features of ImageNet-1K training set requires nearly 20 GiB and 16 ms, respectively, while our CoP and CoRP directly compute the reconstruction error for each new sample with the pre-calculated projection matrix and the mean vector from the training data, resulting in a much higher processing speed and far less storage. The number of RFFs $M$ for CoRP in this experiment is $M=4,096$ ($M\ll N_{\rm tr}$).
In addition, KPCA also outperforms KNN on the CIFAR10 benchmark with improved OoD detection performances, which we leave to Appendix B for more details.
Table 3: Comparisons on the detection performance between the regularized reconstruction error [8] and our CoP and CoRP fused with other OoD scores (MSP, Energy, ReAct and BATS) on each OoD data set (ResNet50 trained on ImageNet-1K). Best average results are highlighted with underlines.
method OoD data sets AVERAGE
iNaturalist SUN Places Textures
@@ 5.2 Comparisons with regularized reconstruction errors
In [8], to alleviate the weak detection performance of PCA reconstruction error $e(\boldsymbol{\hat{z}})$ of Equation 3, the authors proposed to regularize $e(\boldsymbol{\hat{z}})$ by the feature norm $\left\|\boldsymbol{\hat{z}}\right\|_{2}$ and a fusion strategy to boost its detection performance by introducing existing OoD scores. Firstly, the regularized reconstruction error $e_{\rm reg}(\boldsymbol{\hat{z}})$ is calculated in the original $\boldsymbol{z}$-space as : $e_{\rm reg}(\boldsymbol{\hat{z}})=\frac{\|\boldsymbol{U}_{q}\boldsymbol{U}_{q}^{\top}(\boldsymbol{\hat{z}}-\boldsymbol{\mu})-(\boldsymbol{\hat{z}}-\boldsymbol{\mu})\|_{2}}{\left\|\boldsymbol{\hat{z}}\right\|_{2}}.$ Then, the authors claimed that such a regularized version $e_{\rm reg}(\boldsymbol{\hat{z}})$ is still insufficient for OoD detection, and designed a fusion strategy to combine $e_{\rm reg}$ with other existing OoD scores. For example, to fuse $e_{\rm reg}$ with the Energy [5] score, the final scoring function is $S(\boldsymbol{\hat{x}})=(1-e_{\rm reg}(\boldsymbol{\hat{z}}))\cdot S_{\rm Energy}(\boldsymbol{\hat{z}})$.
@@
In this section, we show that our KPCA reconstruction error $e^{\Phi}$ outperforms the regularized PCA reconstruction error $e_{\rm reg}$ under the same fusion framework. Following the settings in [8], for a fair comparison, the fused OoD scores include MSP [15], Energy [5], ReAct [18] and BATS [19]. The detection experiments are executed on the ImageNet-1K benchmark with pre-trained ResNet50 and MobileNet [51] checkpoints from PyTorch [52].
@@
Table 3 presents the comparisons between [8] and ours on the ImageNet-1K benchmark of ResNet50. When fused with MSP, Energy and BATS, both the KPCA-based CoP and CoRP outperform the regularized reconstruction error [8] on almost all the OoD data sets with substantially improved FPR and AUROC values. Specifically, when fused with the ReAct method [18], the CoRP achieves new SOTA OoD detection performance among various prevailing detectors. Experiments on MobileNet also show superior performance of CoP and CoRP, see details in Appendix B.
All these experiment results indicate that an appropriately mapped $\Phi(\boldsymbol{z})$-space benefits the OoD detection, as the non-linearity in $\boldsymbol{z}$-space gets alleviated by the feature mapping $\Phi$. Our work provides 2 viable selections for $\Phi$ with empirical validations, which we hope could attract attentions towards the non-linearity in InD and OoD features for the research community from a kernel perspective.
6 Analytical discussions with KPCA via kernel functions
@@
In CoP and CoRP, KPCA is executed with the covariance matrix of mapped features $\Phi(\boldsymbol{z})$. In contrast, in the classic KPCA [11; 12], such feature mappings $\Phi$ are not explicitly given, and it rather works with a kernel function applied to original features $\boldsymbol{z}$. In this section, we supplement our covariance-based KPCA with its kernel function implementation, including theoretical discussions and empirical comparisons on OoD detection. Our CoP and CoRP are shown to be more effective and efficient than their counterparts that employ kernel functions.
@@
In the classic KPCA, the kernel trick enables projections to the principal subspace via kernel functions without calculating $\Phi$. However, how to map the projections in the principal subspace back to the original $\boldsymbol{z}$-space remains a non-trivial issue, known as the pre-image problem [53], which makes it problematic to calculate reconstructed features via kernel functions. To address this issue, the following Proposition 1 shows a flexible way to directly calculate reconstruction errors without building reconstructed features, so as to apply the kernel trick, shown in the subsequent Proposition 2.
@@
Proposition 1 implies that the reconstruction error equals to the norm of projections in the residual $\boldsymbol{U}^{\Phi}_{p}$-subspace, i.e., the subspace consisting of those principal components that are not kept, see proofs in Appendix D. Accordingly, as typically done in the classic KPCA, we can introduce a kernel function to perform dimension reduction, but to the residual subspace, and then calculate the norms of the reduced features as the reconstruction error, illustrated by Proposition 2.
Given a kernel function $k(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$, we have a kernel matrix $\boldsymbol{K}\in\mathbb{R}^{N_{\rm tr}\times N_{\rm tr}}$ on training data with $\boldsymbol{K}_{i,j}=k(\boldsymbol{z}_{i},\boldsymbol{z}_{j})$, and a vector $\boldsymbol{k}_{\boldsymbol{\hat{z}}}\in\mathbb{R}^{N_{\rm tr}}$ with the $i$-th element $k(\boldsymbol{z}_{i},\boldsymbol{\hat{z}})$ for a new sample $\boldsymbol{\hat{z}}$. Proposition 2 shows how to calculate the KPCA reconstruction error via the kernel function $k$.
@@
The KPCA reconstruction error $e^{k}(\boldsymbol{\hat{z}})$ w.r.t a kernel function $k$ can be calculated as:
@@
where $\boldsymbol{A}\in\mathbb{R}^{N_{\rm tr}\times l}$ includes $l$ eigenvectors of the kernel matrix $\boldsymbol{K}$ w.r.t the top-$l$ smallest eigenvalues.
According to Proposition 2, now CoP and CoRP can be implemented via kernel functions. For CoP, we just directly apply the cosine kernel function $k_{\rm cos}$ on features $\boldsymbol{z}$ to compute the kernel matrix $\boldsymbol{K}$ and the projection matrix $\boldsymbol{A}$, so as to obtain $e^{k}(\boldsymbol{\hat{z}})$ following Equation 9. For CoRP, we should adopt the Gaussian kernel function $k_{\rm gau}$ on the $\ell_{2}$-normalized inputs $\frac{\boldsymbol{z}}{\left\|\boldsymbol{z}\right\|_{2}}$ to calculate $\boldsymbol{K}$, $\boldsymbol{A}$ and $e^{k}(\boldsymbol{\hat{z}})$. Figure 2 shows comparisons on the detection performance between CoP/CoRP and their kernel function implementations.
In Figure 2, the detection performance of KPCA with kernel functions is evaluated by varying the explained variance ratio of the kernel matrix $\boldsymbol{K}$. The larger the explained variance ratio, the smaller the dimension $l$ of $\boldsymbol{A}$. The best detection results achieved by CoP/CoRP are illustrated as the dashed lines. Clearly, regarding the OoD detection performance, reconstruction errors $e^{k}$ calculated by kernel functions are less effective than those calculated explicitly in the mapped $\Phi(\boldsymbol{z})$-space.
@@
Aside from the detection performance, KPCA with kernel functions is far less computationally efficient than CoP/CoRP in two aspects. On the one hand, the time expense of eigendecomposition on the $N_{\rm tr}\times N_{\rm tr}$ kernel matrix $\boldsymbol{K}$ by the former is more expensive than that on the $m\times m$ or $M\times M$ covariance matrix $\boldsymbol{\Sigma}^{\Phi}$ by the latter, since $N_{\rm tr}\gg M$ and $N_{\rm tr}\gg m$. For example, on the ImageNet-1K benchmark with MobileNet, these settings are $N_{\rm tr}=1,281,167$, $M=2560$ and $m=1280$, on which KPCA with kernel functions is actually nearly prohibitive. On the other hand, in the inference stage, KPCA via kernel functions yet requires an $\mathcal{O}(N_{\rm tr})$ time and memory complexity in calculating $\boldsymbol{k}_{\boldsymbol{\hat{z}}}$, as all the training data has to be stored and iterated, which is much higher than the $\mathcal{O}(1)$/$\mathcal{O}(M)$ complexity of our CoP/CoRP.
@@ 7 Conclusion
As PCA reconstruction errors fail to distinguish OoD data from InD data on the penultimate features $\boldsymbol{z}$ of DNNs, kernel PCA is introduced for its non-linearity in the manner of employing explicit feature mappings. To find appropriate kernels that can characterize the non-linear patterns in InD and OoD features, we take a kernel perspective to decouple and analyze key components of an existing KNN detector [7], and thus propose a cosine kernel and a cosine-Gaussian kernel for KPCA. Specifically, two explicit feature mappings $\Phi(\cdot)$ induced from the two kernels are leveraged on original features $\boldsymbol{z}$. For the cosine kernel, its explicit feature mapping can be directly obtained. For the Gaussian kernel, we adopt the celebrated random Fourier features to approximate the Gaussian kernel. The mapped $\Phi(\boldsymbol{z})$-space enables PCA to extract principal components that well separate InD and OoD data, leading to distinguishable reconstruction errors. Extensive empirical results have verified the improved effectiveness and efficiency of the proposed KPCA with new SOTA OoD detection performance. Besides, more in-depth analyzes are drawn on the individual effects of the cosine kernel and the Gaussian kernel, and the involved multiple hyper-parameters. In addition, theoretical discussions and associated experiments are provided to bridge the relationships between our covariance-based KPCA and its kernel function implementation so as to further illustrate the advantages of our method.
@@
One limitation of the KPCA detector is that the two specific kernels are still manually selected with carefully-tuned parameters. It remains a valuable topic in the OoD detection task whether the parameters of kernels could be learned from data according to some optimization objective. For example, deep kernel learning [54] could be considered as an alternative choice so as to pursue stronger kernels that can better characterize InD and OoD with enhanced detection performance by an additional learning step on the features. We hope that the proposed two effective kernels verified empirically in our work could benefit the research community as a solid example for future studies.
Acknowledgments and Disclosure of Funding
This work is jointly supported by National Natural Science Foundation of China (62376153, 62376155), and Shanghai Municipal Science and Technology Research Major Project (2021SHZDZX0102).
Societal impacts
The societal impacts of this work are mainly positive, as it aims at detecting OoD samples in the inference or deployment stage of DNNs, which benefits researches in trustworthy deep learning. Through our work, we hope that new inspirations on the non-linearity in data could be drawn from a kernel perspective so as to highlight the safety issue in real-world machine learning applications.
Appendix A Details of representative OoD detectors
@@
In this section, we elaborate the scoring function $S(\cdot)$ of the OoD detectors included in the comparison experiments of Section 5. Given a well-trained DNN $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{c}$ with inputs $\boldsymbol{x}\in\mathbb{R}^{d}$, the outputs are $c$-dimension logits $f(\boldsymbol{x})\in\mathbb{R}^{c}$ w.r.t $c$ classes. The DNN $f$ learns features $\boldsymbol{z}\in\mathbb{R}^{m}$ of $\boldsymbol{x}$ before the last linear layer, i.e., the penultimate features $\boldsymbol{z}$.
@@
MSP [15] employs the softmax function on the output logits and takes the maximum probability as the scoring function. Given a new sample $\boldsymbol{\hat{x}}\in\mathbb{R}^{d}$, its MSP score is
$S_{\rm MSP}(\boldsymbol{\hat{x}})=\max\left({\rm softmax}(f(\boldsymbol{\hat{x}}))\right).$ (10)
@@
$S_{\rm ODIN}(\boldsymbol{\hat{x}})=\max\left({\rm softmax}(\frac{f(\boldsymbol{\hat{x}}_{\rm a})}{T})\right),$ (11)
@@
Mahala [23] employs the Mahalanobis score to perform OoD detection. The DNN outputs at different layers are modeled as a mixture of multivariate Gaussian distributions, and the Mahalanobis distance is calculated. Then, a linear regressor is trained to achieve a weighted Mahalanobis distance at different layers as the final detection score. To train the linear regressor, the training data and the corresponding adversarial examples are adopted as positive and negative samples, respectively.
$S^{l}_{\rm mahal}(\boldsymbol{\hat{x}})$ $=\max_{i}\ -(f^{l}(\boldsymbol{\hat{x}})-\boldsymbol{\mu}^{l}_{i})^{T}\boldsymbol{\Sigma}_{l}(f^{l}(\boldsymbol{\hat{x}})-\boldsymbol{\mu}_{i}^{l}),$ (12)
@@
Energy [5] uses an energy function on logits as energy is well aligned with input probability densities:
$S_{\rm energy}(\boldsymbol{\hat{x}})=\log\sum_{i=1}^{c}\exp(f_{i}(\boldsymbol{\hat{x}})),$ (13)
where $f_{i}(\boldsymbol{\hat{x}})$ denotes the $i$-th element in the $c$-dimension output logits $f(\boldsymbol{\hat{x}})$.
@@
$\boldsymbol{\bar{z}}=\min\left\{\boldsymbol{z},\alpha\right\},$ (14)
where $\alpha$ is a pre-defined constant. The clipped features $\boldsymbol{\bar{z}}$ then pass through the last linear layer and yield modified logits. Other logits-based OoD methods such as Energy could be applied on the modified logits to produce a detection score.
KNN [7] is a simple but time-consuming and memory-inefficient detector since it performs nearest neighbor search on the $\ell_{2}$-normalized penultimate features between the test sample and all the training samples. The negative of the ($k$-th) shortest $\ell_{2}$ distance is set as the score for a new sample $\boldsymbol{\hat{x}}$.
@@
$f_{\rm DICE}(\boldsymbol{\hat{x}})\triangleq\left(\boldsymbol{M}\odot\boldsymbol{W}\right)^{\top}\boldsymbol{\hat{z}}+\boldsymbol{b}.$ (15)
$\odot$ is the element-wise multiplication, and $\boldsymbol{M}\in\mathbb{R}^{m\times c}$ is a masking matrix whose elements are determined by the element-wise multiplication between the $i$-th column $\boldsymbol{w}_{i}\in\mathbb{R}^{m}$ in $\boldsymbol{W}$ and the penultimate features $\boldsymbol{\hat{z}}$: $\boldsymbol{w}_{i}\odot\boldsymbol{\hat{z}}$. Then, similar as ReAct, logits-based detectors could be executed on the modified logits $f_{\rm DICE}(\boldsymbol{\hat{x}})$ to produce a detection score.